Parth (parthkohli):

\[x^2 + {1 \over x^2} = 7 \Longrightarrow x^3 + {1 \over x^3} = ? \]

OpenStudy (anonymous):

x^4 - 7x^2 + 1 = 0 Let u = x^2 u^2 - 7u + 1 = 0 what are the roots if any?

mathslover (mathslover):

\[\large{(x+\frac{1}{x})^2=x^2+\frac{1}{x^2}+2}\] \[\large{(x+\frac{1}{x})^2=7+2=9}\]

OpenStudy (anonymous):

x3+1/x3 =(a3+b3)=(a+b)(a2-ab+b2)

OpenStudy (anonymous):

x^2 = 7 -1/x^2 multiply by x x^3 = 7x - 1/x

mathslover (mathslover):

from the above we get : \[\large{x+\frac{1}{x}=3}\] right @ParthKohli ?

OpenStudy (anonymous):

just subs in the equation

Parth (parthkohli):

Ah! \(x = 3 - {1 \over x}\)?

mathslover (mathslover):

no dont do that way parth

OpenStudy (anonymous):

\[\large x^3 + \frac{1}{x^3} = (x + \frac{1}{x})^3 - 3(x + \frac{1}{x})\]

mathslover (mathslover):

\[\large{(x+\frac{1}{x})(x^2+\frac{1}{x^2}-1)}\] \[\large{(3)(7-1)}\] this way ..

mathslover (mathslover):

apply the formula of \(\large{x^3+\frac{1}{x^3}}\)

OpenStudy (anonymous):

\[\large x^3 + \frac{1}{x^3} = 3^3 - 3(3)\]

Parth (parthkohli):

Yup. Got it.

mathslover (mathslover):

nice to hear !! best of luck dude

Parth (parthkohli):

Thank you! @mathslover @waterineyes @lgbasallote @higgs @Yahoo! @sai1234

OpenStudy (anonymous):

x3+1/x3 = (x+1/x)(x2+1x2-1) use (x+1/x)2=x2+1x2+2 (x+1/x)2=7+2=9 get (x+1/x) value and substitute in (x+1/x)(x2+1x2-1)